HUBBLE SPACE TELESCOPE DAILY REPORT #1397 PERIOD COVERED: 0000Z (UTC) 05/23/95 - 0000Z (UTC) 05/24/95 Daily Status Report as of 144/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C213.1 as part of a continuum survey and to study the galaxy's morphology. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5980 (The Jet of 3C273) The Planetary Camera was used to image the jet of 3C273. The objective is to study the morphology of the jet at high resolution. The investigators are particularly interested in comparing these data with optical and radio maps to investigate the mechanism for particle acceleration, which occurs all along the jet. The proposal completed with no reported anomalies. 1.5 Completed WF/PC-2 5393 (The Nature of the Host Galaxies of High-Redshift, Radio-Loud Quasars) The WF/PC-2 was used to make observations of galaxy PKS1318+11 which contains a high red-shifted, radio-loud quasar to study its morphology. This data will be compared with low-Z QSOs and other radio galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed FOC/96 5522 (FOC Post-COSTAR Polarization Calibration) The Faint Object Camera (f/96) was used to make polarization calibration observations using standard object NGC-5272 as a source to help find the transmission coefficients. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed FOS/BL 5532 (FOS Dark Monitoring) The Faint Object Spectrograph (Blue Side) was used to monitor the instrument's background (dark) images. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 4 Successful : 4 2.3 Significant HSTARs: None 2.4 Operations Notes: There were no operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-